OpenStudy (anonymous):

What is the solution of the system of equations? You can use any method. y = 3x + 7 y = x – 9 a. (–1, –10) b. (–17, –8) c. (4, 19) d. (–8, –17

OpenStudy (anonymous):

C. Just plug in the x and y values and see if its true or not.

OpenStudy (anonymous):

Here is a example \[-10 = 3(-1) + 7\]\[-10 = -3 + 7\] \[-10 \ne 4\] Therefore A is not the solution. Now you do this for the other three and thats how you find the solution.

OpenStudy (gw2011):

y=3x+7 y=x-9 -3x+y=7 -x+y=-9 Subtract this equation from the above one -2x=16 x=-8 Substitute this value into one of the equations to find y y=x-9 y=-8-9 y=-17 Therefore, the answer is (-8, -17)